# Co-Designing Multimodal Systems for Accessible Asynchronous Dance Instruction

### v3

@@ Abstract.
$c$s: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Accessibility
@@ 1. Introduction
Audio descriptions (AD) can make videos accessible by narrating the visual content that is important to understand the video. While AD is now common for TV shows and movies (Jordan and Oppegaard, 2018), it remains virtually non-existant for user-generated videos on platforms such as YouTube (Liu et al., 2021) and TikTok (Van Daele et al., 2024). Thus, prior work explored manual (YouDescribe, [n. d.]; Liu et al., 2025), semi-automated (Pavel et al., 2020; Liu et al., 2022), and automated (Van Daele et al., 2024; Wang et al., 2021) approaches to adding audio descriptions to videos. But, complex movements that appear in exercise videos remain challenging to describe with verbal narration alone (De Silva et al., 2023) as the AD must convey both what to do (complex body positions, spatial location, and orientation) and how to do it (speed, fluidity, pacing, and expressiveness). Even if the instructions of the video are understandable, it remains challenging for BLV audience members who can not visually compare their own progress to gain feedback (Huh et al., 2025).
@@
We use a participatory design approach to investigate how to make asychronous dance instruction accessible to BLV learners by including BLV learners and dance instructors as equal designers rather than mere consumers of technology. As we aim to maintain the advantages of asynchronous learning from dance tutorial videos available online, we define the solution that our work explores as asychronous dance instruction systems for BLV learners. We select dance as our domain of interest due to the rich complexity in movement mechanics, timing, and multi-movement sequences that makes dance particularly challenging to learn. To co-design asynchronous dance instruction systems, we conduct co-design workshops with participants that have a wide range of relevant expertise in learning, instructing, and communicating around dance: BLV dance learners, dance instructors, technical experts (in AD, sound, and haptics), and participants with multiple types of expertise (e.g., BLV dance instructors). Our workshops aimed to answer the research question: How should we design systems to support asynchronous dance instruction for BLV learners?
@@ 3.1.3. Synchronous, In-Person Workshops for Asynchronous, Remote Dance Instruction Systems
Although our workshops addressed asynchronous and remote dance instruction, we held them in-person to allow for improved collaboration between participants through a shared physical space. The in-person setting allowed participants to establish common ground when learning and teaching one another example dances. Additionally, participants were able to use hands-on prototyping materials when designing their systems. Within their prototypes and during presentations, participants limited system interactions to what would be available in asynchronous contexts. For example, the use of touch was limited to where haptic devices would be placed, or verbal descriptions were limited to what would be provided by the system. Each group was accompanied by a member of the research team to center discussion on asynchronous contexts. Our in-person setup therefore allowed for engaging, collaborative workshops while addressing our problem space of asynchronous dance instruction.
@@ 4.3.3. Tradeoffs Between Modalities
Takeaway: While verbal descriptions are versatile in conveying body mechanics, they are inefficient compared to non-verbal modalities. However, sound and haptic cues are less immediately interpretable than verbal descriptions.
@@ 4.3.5. Customization
Figure 6. System design workflows for groups one to four. Groups used multiple modalities at different stages of the design process including verbal descriptions (purple, or “movement”), use of non-verbal sound (orange, or “sound”, “music”, “metronome”), and use of haptic cues (blue, or “haptic”).
@@ 5. Discussion
We synthesize design implications for potential future systems, reflect on limitations of our workshop, and propose directions for future work.
@@ 5.1. Design Implications
From participants’ system designs, we created 13 design implications (DI1-DI13) for future asynchronous multimodal dance instruction systems (Table 5). These implications address five system goals: convey movement structure, convey timing and expressiveness, instruct in phases, provide feedback, and provide customization.
@@
DI2 V Break individual movements down into verbal descriptions of specific body part motions, body parts in relation to one another, and high-level body positioning and structure. 2, 3, 4, 5, 8 (Rector et al., 2013),(De Silva et al., 2025),(El Raheb et al., 2018)
DI3 V Verbally provide high-level, cross-movement context on dance patterns. 6, 7, 8 (Van Daele et al., 2024)
DI4 H Use haptics to indicate direction of movements 1, 6, 8 (Aggravi et al., 2016),(Hong et al., 2017),(Teng et al., 2025)
Convey timing & expressiveness DI5 V, S Use direct and semantic sound-to-movement mappings and tactile metaphors to convey fluidity, speed, and expressiveness. 1, 2, 3, 4, 5, 7 (De Silva et al., 2025), (Oppici et al., 2020)
DI6 V, S Use non-verbal sound cues and verbal counts to establish tempo. 2, 5, 8 (Keevallik et al., 2024),(Lee et al., 2024)
DI7 H Use haptics to indicate timing of movements. 1, 6, 7, 8 (Aggravi et al., 2016),(Morelli et al., 2010b),(Holland et al., 2018)
Instruct in phases DI8 V, S, H Use learning phases to let users learn movement structure with navigable steps that can be repeated as necessary. 1, 2, 4, 5, 6, 7, 8 (Ulfa and Narawati, 2020),(Hassan et al., 2024), (Fitts and Posner, 1967)
DI9 S, H Use practicing phases to let users practice movements, learn timing, and incorporate dance music. 1, 2, 4, 5, 6, 7, 8 (Zhou et al., 2021)
@@
DI11 S Use sound cues to provide movement structure correctness feedback. 6, 8 (Oh et al., 2013),(Françoise et al., 2016)
@@
Table 5. Design implications for accessible remote dance instruction systems categorized into core system goals, including modalities addressed, participant groups that incorporated each guideline, and related work supporting each guideline. V = verbal, S = sound (non-verbal), H = haptic/tactile. Citations marked with are accessibility-related, are dance-related, and are non-accessibility work using the technique in other tasks.
@@
A common existing approach to dance instruction for all learners is breaking down routines into more manageable pieces, such as into eight-counts (Bläsing, 2015). In our study, all groups similarly segmented routines into individual movements (DI1). While sighted learners might rely on visual demonstrations to learn each segmented movement (Bläsing et al., 2018; Bläsing, 2015), BLV learners in our workshops further incorporated a second level of segmentation with verbal descriptions of individual body parts and what they were doing (DI2). As BLV learners can not directly preview the video or instructor, providing high-level context to frame (e.g., “this dance works in opposites; every time you step forward and vice versa”) the dance is also valuable (DI3).
In our study, participants typically used their own vocabularies to name and describe movements, occassionally using an existing movement term (e.g., triple-step, Plié). Dance styles such as Ballet, Cha-cha, and Hip-Hop contain standardized, named movements or sequences (Atã and Queiroz, 2019; Bennett, 2008) that can be used in future systems. Similarly, Laban Movement Analysis (LMA)  (Groff, 1995) explicitly categorizes movements based on what parts are moving, orientation and direction of these parts, overall body shape, and dynamics such as weight and flow. To increase consistency between dances, instructors or future systems may use such existing dance vocabularies or frameworks to name movements.
Furthermore, BLV participants raised common stategies from their lived experience for conveying orientation and spatial information such as using the clock to convey orientation (Tabb, 2025). Thus, systems could tailor their descriptions using best practices from orientation and mobility training or BLV learners’ own preferred terminology.
@@ 5.1.2. Convey Timing and Expresssiveness
Many dance styles contain temporal (speed-based) and expressive (feeling-based) qualities that are critical for dance instruction (Blom and Chaplin, 2004). Among our participants, non-verbal sound cues and metaphors proved to be a popular way to convey such aspects that are difficult to objectively describe. Existing movement analysis work (Frid et al., 2018; Landry and Jeon, 2020) shows that sound effectively conveys temporal and expressive movement qualities. Our BLV participants emphasized having sound-to-movement mappings in their dance instruction systems as an efficient way of grasping these movement qualities that sighted learners are able to perceive visually — for example, the duration of the sound cue matching movement duration. Prior work on accessible dance (De Silva et al., 2023; Bläsing and Zimmermann, 2021) additionally highlights the importance of using metaphors to verbally convey these nuanced movement qualities to BLV individuals.
The use of non-verbal sound beats and verbal counts to indicate rhythm (DI6) is also prevalent in existing dance literature (Keevallik et al., 2024; Lee et al., 2024). Our workshops showed that these techniques provide additional grounding for BLV participants who rely more heavily on auditory information in place of visual demonstrations.
@@
Accessible remote dance instruction systems should teach in phases (DI8-DI9). Our results conclude that early instructional phases should prioritize detail in vocabulary definitions, verbal descriptions, and specific feedback as users internalize movement structure. Following this, users are able to build practice through repetition and the emphasis on non-verbal cues to establish timing and expressive qualities. Users are then able to explore how the routine interacts with music by overlaying the music on non-verbal cues, before eventually removing all cues. BLV participants emphasized repetition as important for developing body awareness and internalize choreography (DI8). Systems should thus allow for navigable steps that users can revisit and repeat while putting routines together. These instructional phases reflect existing methods for motor skill learning, where learners gradually increase autonomy with repetition and practice (Hassan et al., 2024; Fitts and Posner, 1967; Ulfa and Narawati, 2020). Our work shows that such repetition and practice is particularly important in accessible dance instruction systems to allow BLV learners to form accurate mental models of spatial information that they are not able to visually perceive.
@@ 5.1.4. Provide Feedback
Our results showed that BLV learners benefit from immediate feedback to confirm orientation, spatial positioning, and overall confidence in their movements. Thus, we derived multimodal feedback techniques (DI10-DI12). Movement correction and feedback is widely used in existing dance practice. However, the ambiguity of verbal descriptions is higher for BLV learners (De Silva et al., 2025; Bläsing and Zimmermann, 2021), while sighted learners are able to implicitly receive feedback through visual comparison of their movements (Bläsing et al., 2018). BLV participants thus emphasized having clarity in their movement goals and frequent feedback on their movements. Future accessible dance instruction systems can incorporate this model by providing clear descriptions of movement goals (DI2), multimodal feedback on movement correctness (DI10-DI12), and incremental feedback to correct positioning and orientation (DI12).
@@ 5.1.5. Provide Customization
As outlined by participants and prior work (Aflatoony and Lee, 2020; Jones et al., 2024; Kane et al., 2014), customizable technologies improve the technology experience of people with disabilities. Our results showed clear differences across learners: some valued detailed auditory explanations, while others preferred tactile or haptic cues that allowed them to physically establish orientation and weight. Users additionally had varied level of detail required in descriptions. For example, participants B5a, B6, B7b and B8 had prior dancing experience, allowing them to better follow the dance to the point where they could teach other participants. Comparatively, BLV participants with no experience required more explicit instructions and greater detail to learn the same dances. Asynchronous dance instruction systems should thus allow for customizable experiences that acccount for prior knowledge of dance style and skill level (DI13). Our workshops additionally revealed the need for system differences across dance styles (DI14), which is discussed in section 5.2.
@@ 5.2.1. Example Accessible Asynchronous Dance Instruction Systems
In a preprocessing phase, the system takes a video of a choreographed routine and segments it into a set of individual movements (DI1) (e.g., using a Temporal Convolutional Network to segment distinct dance movements as in prior work (Endo et al., 2024)). Next, the system provides detailed text descriptions of each movement (DI2) (e.g., with a motion-specific vision language model (Chen et al., 2024)). For each segmented movement, the system also creates sound-to-movement mappings (DI5). Direct mappings are created by using pose detection models such as MediaPipe (Sengar et al., 2024; Lugaresi et al., 2019) and AlphaPose (Fang et al., 2022) to extract joint data and sonify them (e.g., mapping pitch to height or volume to speed). To create semantic mappings, the system may first generate expressive descriptions of each movement (Yazdian et al., 2025) and then create expressive sound cues from these descriptions with a text-to-audio model (Huang et al., 2023). During system setup, users may include information on prior dance experience to provide a customized learning experience (DI13) — e.g., the system may substitute detailed low-level descriptions for familiar dance terms for if the participant is familiar with the dance style.
@@
Participants can then use the system by freely navigating via screen reader or voice command between different movements to learn and practice each motion by listening to the verbal and sound cues (DI8). Once a user is comfortable with steps, they may practice them at full speed and on top of existing dance music during practice phases (DI9). Throughout these phases, the system will further collect a video stream of the user’s attempts for each step and provide verbal corrections (DI12) and positive sound cues once a user performs a step correctly (DI11).
Virtual Dance Partner System: Informed by G1, G6 and G8 designing systems for Salsa and Cha-cha, we envision a virtual dance partner for teaching steps from social dance videos that often demonstrate repeating standardized steps. Similar to G8, this system will teach a set of standardized steps (e.g., rock-step, triple-step) (DI1) using verbal descriptions (D12) combined with a pair of haptic probes on the learner’s back that will vary haptic vibration intensity and location corresponding to weight distribution and direction of foot placement (DI4) (e.g., similar to body worn haptics for acrobatics (Popp et al., 2023)). The system will include wrist-worn haptics (Hong et al., 2017) that vibrate to cue movements (DI4, DI7) (i.e., left wrist vibrating means to step left, a double vibration means to turn). Once learning the named movements, users are able to practice these steps with haptic cues (DI9).

### v4

# arXiv:2511.09658v4[cs.HC] 27 Feb 2026

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791376isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Human computer interaction (HCI)$c$s: Human-centered computing Accessibility
@@ 1. Introduction
Audio descriptions (AD) can make videos accessible by narrating the visual content that is important to understand the video. While AD is now common for TV shows and movies (Jordan and Oppegaard, 2018), it remains virtually non-existant for user-generated videos on platforms such as YouTube (Liu et al., 2021) and TikTok (Van Daele et al., 2024). Thus, prior work explored manual (YouDescribe, [n. d.]; Wang et al., 2025), semi-automated (Pavel et al., 2020; Liu et al., 2022), and automated (Van Daele et al., 2024; Wang et al., 2021) approaches to adding audio descriptions to videos. But, complex movements that appear in exercise videos remain challenging to describe with verbal narration alone (De Silva et al., 2023) as the AD must convey both what to do (complex body positions, spatial location, and orientation) and how to do it (speed, fluidity, pacing, and expressiveness). Even if the instructions of the video are understandable, it remains challenging for BLV audience members who can not visually compare their own progress to gain feedback (Huh et al., 2025).
@@
We use a participatory design approach to investigate how to make asynchronous dance instruction accessible to BLV learners by including BLV learners and dance instructors as equal designers rather than mere consumers of technology. As we aim to maintain the advantages of asynchronous learning from dance tutorial videos available online, we define the solution that our work explores as asynchronous dance instruction systems for BLV learners. We select dance as our domain of interest due to the rich complexity in movement mechanics, timing, and multi-movement sequences that makes dance particularly challenging to learn. To co-design asynchronous dance instruction systems, we conduct co-design workshops with participants that have a wide range of relevant expertise in learning, instructing, and communicating around dance: BLV dance learners, dance instructors, technical experts (in AD, sound, and haptics), and participants with multiple types of expertise (e.g., BLV dance instructors). Our workshops aimed to answer the research question: How should we design systems to support asynchronous dance instruction for BLV learners?
@@ 3.1.3. Synchronous, In-Person Workshops for Asynchronous, Remote Dance Instruction Systems
Although our workshops addressed asynchronous dance instruction, we held them in-person to allow for improved collaboration between participants through a shared physical space. The in-person setting allowed participants to establish common ground when learning and teaching one another example dances. Additionally, participants were able to use hands-on prototyping materials when designing their systems. During prototyping and presentations, participants limited system interactions to what would be available in asynchronous contexts. For example, the use of touch was limited to where haptic devices would be placed, or verbal descriptions were limited to what would be provided by the system. Each group was accompanied by a member of the research team to center discussion on asynchronous contexts. Our in-person setup therefore allowed for engaging, collaborative workshops while addressing our problem space of asynchronous dance instruction.
@@ 4.3.3. Tradeoffs Between Modalities
Takeaway: While verbal descriptions are versatile in conveying body mechanics, they are inefficient compared to non-verbal modalities and leave room for ambiguity. However, sound and haptic cues should be simplified and provided in moderation to reduce cognitive load.
@@ 4.3.5. Customization
Figure 6. System design workflows for groups one to four. Groups used multiple modalities at different instructional stages, including verbal descriptions (purple, or “movement”), non-verbal sound (orange, or “sound”, “music”, “metronome”), and haptic cues (blue, or “haptic”).
@@ 5. Discussion
We synthesize design implications for potential future systems, reflect on limitations of our workshops, and propose directions for future work.
@@ 5.1. Design Implications
From participants’ system designs, we created 14 design implications (DI1-DI14) for future asynchronous multimodal dance instruction systems (Table 5). These implications address five system goals: convey movement structure, convey timing and expressiveness, instruct in phases, provide feedback, and provide customization.
@@
DI2 V Break individual movements down into verbal descriptions of specific body part motions, body parts in relation to one another, and high-level body positioning. 2, 3, 4, 5, 8 (Rector et al., 2013),(De Silva et al., 2025),(El Raheb et al., 2018), (Groff, 1995)
DI3 V Verbally provide high-level context on dance patterns. 6, 7, 8 (Van Daele et al., 2024)
DI4 H Use haptics to indicate movement direction 1, 6, 8 (Aggravi et al., 2016),(Hong et al., 2017),(Teng et al., 2025), (Popp et al., 2023)
Convey timing & expressiveness DI5 V, S Use sound-to-movement mappings and tactile metaphors to convey temporal and expressive qualities. 1, 2, 3, 4, 5, 7 (De Silva et al., 2025), (Oppici et al., 2020)
DI6 V, S Use non-verbal beats and verbal counts to establish rhythm. 2, 5, 8 (Keevallik et al., 2024),(Lee et al., 2024)
DI7 H Use haptics to indicate movement timing. 1, 6, 7, 8 (Aggravi et al., 2016),(Morelli et al., 2010b),(Holland et al., 2018), (Popp et al., 2023)
Instruct in phases DI8 V, S, H Use learning phases to let users learn movement structure and multimodal cue definitions. 1, 2, 4, 5, 6, 7, 8 (Ulfa and Narawati, 2020),(Hassan et al., 2024)
DI9 S, H Use practicing phases once users have learned movements to allow repetition, practice timing, and incorporate dance music. 1, 2, 4, 5, 6, 7, 8 (Zhou et al., 2021),(Fitts and Posner, 1967)
@@
DI11 S Use sound cues to provide correctness feedback. 6, 8 (Oh et al., 2013),(Françoise et al., 2016)
@@
Table 5. Design implications for accessible asynchronous dance instruction systems categorized into core system goals, including modalities addressed, participant groups that incorporated each implication, and related work supporting each implication. V = verbal, S = sound (non-verbal), H = haptic/tactile. Citations marked with are accessibility-related, are dance-related, and are non-accessibility work using the technique in other tasks.
@@
An existing dance instruction approach is breaking down routines into manageable pieces, such as eight-counts (Bläsing, 2015). Without access to visual demonstrations of these pieces, BLV learners relied on a multimodal movement vocabulary of named movements (DI1). Future systems may draw from past work using Temporal Convolutional Networks (TCNs) (Endo et al., 2024) and vision-language models (VLMs) (Huh et al., 2025) to segment videos into discrete steps.
While sighted learners might rely on visual demonstrations to learn each segmented movement (Bläsing et al., 2018; Bläsing, 2015), BLV participants incorporated a second level of segmentation with granular verbal descriptions of individual body parts (DI2) and haptic cues to indicate movement direction (DI4). Future systems may explore generating detailed descriptions through established VLM-based movement understanding techniques (Chen et al., 2024) and motion-to-text models (Jiang et al., 2024). Such systems may draw from frameworks such as Laban Movement Analysis  (Groff, 1995), which categorizes movements based on moving body parts, orientation/direction of these parts, and overall body shape. Dance styles such as Ballet, Cha-cha, and Hip-Hop contain standardized, named movements (Atã and Queiroz, 2019; Bennett, 2008) which can be incorporated into systems for greater consistency. Wearable haptic bracelets can additionally indicate movement directions (Hong et al., 2017) (e.g., the front of your ankle vibrating meaning to move your foot forward).
As BLV learners cannot visually preview routines, providing high-level context (e.g., “this dance works in opposites; every time you step forward, step backwards”) is valuable (DI3). Existing tools for accessible video descriptions (Van Daele et al., 2024) provide high-level summaries of video content, which may be extended to dance routines.
@@ 5.1.2. Convey Timing and Expresssiveness
Many dances contain temporal (time-based) and expressive (feeling-based) qualities (Smith-Autard, 2014). Participants used non-verbal sound cues and metaphors to convey such aspects that are difficult to describe concisely (DI5)e.g., the duration of the sound matching movement duration. Future systems may draw from existing work using sound to convey movement emotion and fluidity (Frid et al., 2018; Landry and Jeon, 2020). Prior work on accessible dance (De Silva et al., 2023; Bläsing and Zimmermann, 2021) additionally highlights using metaphors to convey movement qualities. Our work preliminarily found that tactile metaphors were most commonly used for accessible descriptions. Future work may involve more formal evaluations of effective metaphor types for describing complex movement.
Sound beats and verbal counts to indicate rhythm (DI6) is common in existing instruction (Keevallik et al., 2024; Lee et al., 2024). Our workshops showed that these techniques provide crucial grounding for BLV participants who rely on auditory information in place of visual demonstrations. Future dance instruction systems may therefore include a metronome beat and/or verbal counts as part of their movement descriptions. Participants also found haptics useful for cueing when to move (DI7). Future systems can incorporate wearable haptics with vibrations timed with verbal instructions. However, such systems should consider limiting the number of haptic devices to reduce cognitive load.
@@
Existing methods for motor skill learning involve gradually increasing movement autonomy with repetition and practice (Hassan et al., 2024; Fitts and Posner, 1967; Ulfa and Narawati, 2020). Our work shows that such structured learning, repetition, and practice is particularly important for allowing BLV learners to internalize spatial information that they cannot visually perceive. Accessible asynchronous dance instruction systems may include early instructional phases that prioritize detail in movement definitions, verbal descriptions, and specific feedback as users internalize movement structure (DI8). Learners can then practice with non-verbal cues to build muscle memory and establish timing and expressive qualities (DI9). BLV participants emphasized repetition as important for developing body awareness and internalizing choreography. Future systems should thus allow for repeatable segments that users can practice (DI9).
@@ 5.1.4. Provide Feedback
Feedback is widely used in dance instruction. However, the ambiguity of verbal descriptions is higher for BLV learners (De Silva et al., 2025; Bläsing and Zimmermann, 2021), while sighted learners can implicitly receive feedback through visual comparison (Bläsing et al., 2018). BLV participants thus emphasized frequent, multimodal feedback to confirm orientation and structure and increase confidence in their movements.
Participants found physical elements helpful for providing spatial/orientation feedback (DI10). Systems may incorporate textured or haptic mats with markings to indicate orientation. Participants also found sound cues (DI11), such as a bell sound when a movement is done correctly, useful for correctness feedback and encouragement. For complex movements involving multiple body parts, participants emphasized incremental verbal feedback (DI12). Future systems can use pose models (Fang et al., 2022) to compare learner attempts and reference videos and suggest corrections (Rector et al., 2013) or provide sound feedback.
@@ 5.1.5. Provide Customization
As outlined by participants and prior work (Aflatoony and Lee, 2020; Jones et al., 2024; Kane et al., 2014), customizable technologies improve the experience of people with disabilities. Our results showed clear differences across learners: some valued detailed auditory explanations, while others preferred tactile or haptic cues that allowed them to physically establish orientation and weight. Users additionally required varied levels of description detail. For example, participants B5a, B6, B7b and B8 had prior dancing experience, allowing them to better follow the dance to the point where they could teach other participants. Comparatively, BLV participants with no experience required more detailed instructions to learn the same dances. Asynchronous dance instruction systems should thus allow for customizable experiences that acccount for prior knowledge of dance style and skill level (DI13). Our workshops additionally revealed system differences across dance styles (DI14), discussed in section 5.2.
@@ 5.2.1. Example Accessible Asynchronous Dance Instruction Systems
In a preprocessing phase, the system takes a video of a choreographed routine and segments it into a set of individual movements (DI1) (e.g., using a Temporal Convolutional Network to segment distinct dance movements as in prior work (Endo et al., 2024)). Next, the system provides detailed text descriptions of each movement (DI2) (e.g., with a motion-specific vision language model (Chen et al., 2024)). For each segmented movement, the system also creates sound-to-movement mappings (DI5). Direct mappings are created by using pose detection models such as MediaPipe (Sengar et al., 2024; Lugaresi et al., 2019) and AlphaPose (Fang et al., 2022) to extract joint data and sonify them (e.g., mapping pitch to height or volume to speed). To create semantic mappings, the system may first generate expressive descriptions of each movement (Yazdian et al., 2025) and then create expressive sound cues from these descriptions with a text-to-audio model (Huang et al., 2023). During system setup, users may include information on prior dance experience to provide a customized learning experience (DI13)e.g., the system may substitute detailed low-level descriptions for familiar dance terms for if the participant is familiar with the dance style.
@@
Participants can then use the system by freely navigating via screen reader or voice command between different movements to learn each motion by listening to the verbal and sound cues (DI8). Once a user is comfortable with steps, they may practice them at full speed with non-verbal cues and existing dance music during practice phases (DI9). Throughout these phases, the system will further collect a video stream of the user’s attempts for each step and provide verbal corrections (DI12) and positive sound cues once a user performs a step correctly (DI11).
Virtual Dance Partner System: Informed by G1, G6 and G8 designing systems for Salsa and Cha-cha, we envision a virtual dance partner for teaching steps from social dance videos that often demonstrate repeating standardized steps. During learning phases (DI8), this system will cover a set of pre-programmed, standardized steps (e.g., rock-step, triple-step) (DI1) using verbal descriptions (D12) combined with a set of back-worn haptic devices (DI4). Each step pattern will be individually defined with verbal descriptions of feet placement and pre-programmed haptic demonstrations of the step. We envision a total of four vibrotactile motors positioned horizontally across the learner’s back, where vibrations indicate foot contact with the ground. Having four motors will allow the system to demonstrate steps where feet cross over one another, such as in the triple-step. Haptic vibration intensity will correspond to weight distribution of the steps (e.g., a light vibration for a toe tap, a heavier vibration for weight shifts).
The system will include wrist-worn haptics (Hong et al., 2017) that vibrate to cue movements and indicate direction (DI4, DI7) (e.g., left wrist vibrating means to step left, a double vibration means to turn). Once learning the named movements, users are able to practice these steps with the wrist-based haptic cues, before eventually removing all cues (DI9). Users are able to adjust vibration intensity for the haptic devices based on their preference, or decide to remove any set of haptics (e.g., no back-worn haptics during learning) (DI13).